# Periodic Asynchrony: An Effective Method for Accelerating On-Policy Reinforcement Learning

### v1

@@ 4.1 Unified Tri-Model Architecture
For the third step, mainstream reinforcement learning systems typically require the computation of three types of logits in the forward pass: policy, old policy, and reference. These models share the same architecture, differing only in their weights. In this work, the training phase is implemented using a Megatron-like architecture, where a complete policy model is distributed across multiple computational units according to tensor parallelism and pipeline parallelism scales. During the forward computation of the old policy and reference models, each sub-network managed by the policy model is replicated twice across all computational units—once for computing the old policy logits and once for computing the reference logits. The reference network retains the original model weights, while the old policy network holds weights delayed by one training step relative to the policy model. The overall architecture of the reinforcement learning system constructed in this work is illustrated in Figure1.
@@ 4.2 Shared-Prompt Attention
    The position IDs need to be rearranged. For the second and subsequent responses, their starting positions should immediately follow the prompt, maintaining continuity with the prompt tokens.
@@
4.3 Asynchronous On-Policy Reinforcement Learning
@@
Algorithm 1 Periodic Asynchronous On-Policy Reinforcement Learning Algorithm
@@
Figure 4: Comparison of training steps in synchronous and asynchronous systems.
@@
Then, all samples in this batch are trained sequentially based on the order in which they enter the queue. Therefore, theoretically, if the training prompts in a batch cannot be processed by a single inference batch, the asynchronous system can achieve more than twice the speed of the synchronous system.
@@ 6 Experiments
The main objective is to experimentally validate the performance between periodic asynchronous and synchronous training, as well as to compare the proposed periodic asynchronous system with other mainstream synchronous reinforcement learning frameworks on the same platform, while also evaluating the scalability of our periodic asynchronous system.
The primary evaluation metric is the end-to-end training throughput, such as the average number of tokens trained per second per device (TPSPD). This is calculated as the total number of tokens processed over several steps divided by the total elapsed time and then divided by the number of devices used.
Since no modifications were made to the reinforcement learning algorithm, accuracy metrics are reported only for reference, primarily to verify the correctness of algorithm reproduction.
@@ 6.2 Comparison between Asynchronous and Synchronous Training
Sync (our) 100 0.746 827.306M 143.678 16 99.966
Async (our) 100 0.758 917.609M 82.861 16 192.259
@@ 6.3 Asynchronous Scalability Experiments
We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group. Under an inter-node bandwidth of 100 Gb/s, the results are shown in Table 3.
@@ 7 Conclusion
This paper presents an asynchronous on-policy reinforcement learning implementation that maintains the same training stability as synchronous RL methods. The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated. Prompt-level asynchronous parallelism maximizes the overlap between inference and training. During training, a three-mode (policy, old policy, reference) unified and identically distributed architecture is introduced, along with a group-mask attention mechanism that provides clear advantages when prompts are long and responses are short.
Experiments show that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform. The results also support the equivalence in training effectiveness between the asynchronous and synchronous methods. Furthermore, the decoupling of training and inference allows both to be scaled independently and flexibly, achieving optimal balance and maximizing the speed-to-device ratio. The decoupling of training architecture and algorithm also implies that the proposed framework can be applied to other on-policy reinforcement learning algorithms without any modifications for compatibility with asynchronous training architectures.

### v2

# Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning

@@ 4.1 Unified Tri-Model Architecture
For the third step, mainstream reinforcement learning systems typically require the computation of three types of logits in the forward pass: policy, old policy, and reference. These models share the same architecture, differing only in their weights. In this work, the training phase is implemented using a Megatron-like architecture, where a complete policy model is distributed across multiple computational units according to tensor parallelism and pipeline parallelism scales. During the forward computation of the old policy and reference models, each sub-network managed by the policy model is replicated twice across all computational units—once for computing the old policy logits and once for computing the reference logits. The reference network retains the original model weights, while the old policy network holds weights delayed by one training step relative to the policy model.
@@
The overall architecture of the reinforcement learning system constructed in this work is illustrated in Figure 1. The three types of models share an identical architecture and are distributed in the same parallel layout. This design allows their logits to be computed simultaneously without requiring any additional topology for the reference model. The inference stage, in contrast, is handled by an independent inference engine such as vLLM. Before training each batch, the weights of the policy model are synchronized to all inference engines. The inference engines then receive prompts from the dataset, generate responses, and return them for scoring. Each scored response is concatenated with its corresponding prompt to form a training sample, which is fed into all three models for forward computation. In a single micro-step, we obtain the logits from the policy, old policy, and reference models simultaneously. After all samples in a batch complete their forward passes—and before the policy model’s weights are updated—its current weights are moved into the old policy model’s state_dict.
@@ 4.2 Shared-Prompt Attention
    The position indices need to be rearranged. For the second and subsequent responses, their starting positions should immediately follow the prompt, maintaining continuity with the prompt tokens.
@@
4.3 Periodic Asynchronous Reinforcement Learning
@@
Algorithm 1 Periodic Asynchronous Reinforcement Learning Algorithm
@@
Figure 4: Comparison of training steps in synchronous and our asynchronous systems.
@@
After that, all samples in the batch are trained sequentially in the order in which they enter the queue. It is clear that when inference speed and training speed are well balanced, the sum of their waiting times is minimized, resulting in optimal end-to-end training throughput on a given hardware configuration. The decoupled inference–training architecture naturally enables such balancing by allowing independent configuration of inference and training instances.
Therefore, in theory, when a training batch cannot be processed within a single inference batch, the asynchronous system can achieve more than twice the startup speed of the synchronous system.
@@ 6 Experiments
The main objective is to experimentally validate the performance between periodic asynchronous and synchronous training, as well as to compare the proposed periodic asynchronous system with other mainstream synchronous reinforcement learning frameworks on the same platform, while also evaluating the scalability of our periodic asynchronous system. The primary evaluation metric is the end-to-end training throughput, such as the average number of tokens trained per second per device (TPSPD). This is calculated as the total number of tokens processed over several steps divided by the total elapsed time and then divided by the number of devices used. Since no modifications were made to the reinforcement learning algorithm, accuracy metrics are reported only for reference, primarily to verify the correctness of algorithm reproduction.
The experiments in this study were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight NPUs, each providing 64 GB of memory. The intra-node interconnect bandwidth is 196 GB/s, while the inter-node bandwidth is 100 Gb/s.
@@ 6.2 Comparison between Asynchronous and Synchronous Training
Sync (ours) 100 0.746 827.306M 143.678 16 99.966
Async (ours) 100 0.758 917.609M 82.861 16 192.259
@@ 6.3 Asynchronous Scalability Experiments
We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group. The results are shown in Table 3.
@@ 7 Conclusion
This paper presents an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods. The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated. Prompt-level asynchronous parallelism maximizes the overlap between inference and training. During training, a three-mode (policy, old policy, reference) unified and identically distributed architecture is introduced, along with a group-mask attention mechanism that provides clear advantages when prompts are long and responses are short. Experiments show that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform. The results also support the equivalence in training effectiveness between the asynchronous and synchronous methods. Furthermore, the decoupling of training and inference allows both to be scaled independently and flexibly, achieving optimal balance and maximizing the speed-to-device ratio. The decoupling of training architecture and algorithm also implies that the proposed framework can be applied to other on-policy reinforcement learning algorithms without any modifications for compatibility with asynchronous training architectures.